# State Combinatorial Generalization In Decision Making With Conditional Diffusion Models

### v1

@@ D.8.1 Experiment Details
D.8.2 Model Size
We include the model size for different algorithms below in Table 5. To eliminate the concern for performance gain due to model size, we include the performance of a large BC model and PPO that has roughly the same number of parameters as the conditioned diffusion model.
BC PPO Diffusion Large BC Large PPO
Model size 0.30 MB 0.60 MB 54.19 MB 55.65 MB 111.29MB (Policy:55.64+Value:55.64)
Number of parameters 75013 148998 13546370 13912325 27823622 (Policy:13911040 + Value:13911040)
OOD reward 7.50 (0.03) 8.19 (0.16) 8.81 (0.2) 7.71 (0.3) 8.43 (0.19)
OOD crashes 37.7 (0.5) 36.0 (5.7) 20.0 (2.5) 37.3 (4.0) 31.67 (1.89)
@@ D.8.2 Model Size
D.8.3 Reliable Conditioning
We demonstrate the importance of having reliable information of base element composition, we compare the performance of the conditioned diffusion model given random and ground truth conditionings.
Ground Truth Random
Number of Crashes 19.67 (2.49) 24.33 (0.47)
Reward 8.81(0.2) 8.1 (0.09)
Table 6: Performance of conditioned diffusion model given ground truth and random conditionings.
D.9 StarCraft
D.9.1 Experiment Details
@@
D.9.2 Dataset Initial State Distribution
The probability of generating each unit type in SMACv2 is imbalanced. Specifically, the probability for Stalker, Zealot, and Colossus is 0.45, 0.45, and 0.1 respectively. The initial state distribution of training trajectories collected by MAPPO for random combination is shown below in Figure 13(a) and 13(b). Since we only keep the successful trajectories and use them as expert data, the distribution depends on the generation probability and MAPPO success rate for different team combinations. A total number of 240000 trajectories were used to train the diffusion model. Since diffusion is trained on local observations and actions of all MAPPO actors, the total number of training samples is 5*240000 for 5v5 and 3*240000 for 3v3.
@@ D.9.2 Dataset Initial State Distribution
D.9.3 Detailed Results on SMACv2
Table  10 and 11 show the detailed performance of different algorithms in the 3v3 and 5v5 scenarios, respectively.
@@ 2 PPO + 1 Rand 3 PPO BC 2 PPO + 1 Diffusion
Table 10: Success rate of each agent in 100 rounds. The first two rows correspond to the simple setting of generalization to states with different support and the last two rows correspond to non-overlapping support. Numbers in the parenthesis represent the standard error over 3 seeds. The best performing method is labeled bold. The 2 PPO + 1 Rand column shows the effect of replacing one PPO trained agent with a random agent as a baseline for comparison against the 2 PPO + 1 Diffusion case.
@@ 4 PPO + 1 Rand 5 PPO BC 4 PPO + 1 Diffusion
Table 11: Success rate of each agent in 100 rounds. The first two rows correspond to the simple setting of generalization to states with different support and the last two rows correspond to non-overlapping support. Numbers in the parenthesis represent the standard error over 3 seeds. The best performing method is labeled bold. The 4 PPO + 1 Rand column shows the effect of replacing one PPO trained agent with a random agent as a baseline for comparison against the 4 PPO + 1 Diffusion case.
D.9.4 Detailed Results for Ablation
@@
$ABC\to ABC$ (ID) 0.53$\pm$0.04 0.59$\pm$0.03 0.66$\pm$0.01
$ABC\to AAA$ (Simple) 0.50$\pm$0.03 0.50$\pm$0.01 0.56$\pm$0.02
$AAA\to AAA$ (ID) 0.47$\pm$0.08 0.55$\pm$0.03 0.58$\pm$0.04
$AAA\to ABC$ (Hard) 0.27$\pm$0.03 0.32$\pm$0.04 0.30$\pm$0.04
D.9.5 2v2
The success rates for StarCraft 2v2 are shown below in Table 14. We can see that out-of-combination cases did not cause the performance to drop drastically for MAPPO. This is because the number of combinations in 2v2 is very limited (e.g. $aa,bb,ab$), and if one agent dies, MAPPO has encountered scenarios of playing with each unit type individually, therefore falling back to in distribution state again. This scenario also exists for 5v5 and 3v3 but only at the end of each game when only one agent is left.
@@
$ABC\to ABC$ (ID) 0.54$\pm$0.02 0.06$\pm$0.02 0.62$\pm$0.05 0.43$\pm$0.04 0.56$\pm$0.03 0.57$\pm$0.067
$ABC\to AAA$ (Simple) 0.47$\pm$0.02 0.02$\pm$0.01 0.57$\pm$0.02 0.44$\pm$0.02 0.55$\pm$0.06 0.49$\pm$0.06
$AAA\to AAA$ (ID) 0.57$\pm$0.01 0.01$\pm$0.01 0.64$\pm$0 0.38$\pm$0.02 0.63$\pm$0.04 0.63$\pm$0.02
$AAA\to ABC$ (Hard) 0.4$\pm$0.03 0.04$\pm$0.02 0.44$\pm$0.02 0.29$\pm$0.02 0.43$\pm$0.08 0.41$\pm$0.06
D.9.6 More Rendering of States Predicted by the Diffusion model
@@
D.10 Model Runtime and GPU Memory
We include the training time and GPU memory used for the conditioned diffusion model below in Table 15 and 16.
@@ D.10 Model Runtime and GPU Memory
D.11 Parameter Comparison with Concatenation or Attention
We demonstrate the number of parameters in attention-based conditioning and concatenation-based conditioning to eliminate the concern regarding performance gain due to more parameters. Attention or concatenation has roughly the same number of parameters as the attention module is convolutional layers and concatenation increases the parameters of conditioning layer.
Attention Concatenation
Model size 617.06 MB 619.64 MB
Parameters 154264085 154911187
Table 17: Number of parameters in attention-based conditioning and concatenation-based conditioning.
D.12 Substituting More MAPPO Agents with Diffusion Agents

### v2

@@ C.7.1 Experiment Details
C.7.2 Model Size
We include the model size for different algorithms below in Table 5. To eliminate the concern for performance gain due to model size, we include the performance of a large BC model that has roughly the same number of parameters as the conditioned diffusion model.
BC PPO Diffusion Large BC
Model size 0.30 MB 0.60 MB 54.19 MB 55.65 MB
Number of parameters 75013 148998 13546370 13912325
OOD reward 7.50 (0.03) 8.19 (0.16) 8.81 (0.2) 7.71 (0.3)
OOD crashes 37.7 (0.5) 36.0 (5.7) 20.0 (2.5) 37.3 (4.0)
@@ C.7.2 Model Size
C.8 StarCraft
C.8.1 Experiment Details
@@
C.8.2 Dataset Initial State Distribution
The probability of generating each unit type in SMACv2 is imbalanced. Specifically, the probability for Stalker, Zealot, and Colossus is 0.45, 0.45, and 0.1 respectively. The initial state distribution of training trajectories collected by MAPPO for random combination is shown below in Figure 12(a) and 12(b). Since we only keep the successful trajectories and use them as expert data, the distribution depends on the generation probability and MAPPO success rate for different team combinations. A total number of 240000 trajectories were used to train the diffusion model. Since diffusion is trained on local observations and actions of all MAPPO actors, the total number of training samples is 5*240000 for 5v5 and 3*240000 for 3v3.
@@ C.8.2 Dataset Initial State Distribution
C.8.3 Detailed Results on SMACv2
Table  9 and 10 show the detailed performance of different algorithms in the 3v3 and 5v5 scenarios, respectively.
@@ 2 PPO + 1 Rand 3 PPO BC 2 PPO + 1 Diffusion
Table 9: Success rate of each agent in 100 rounds. The first two rows correspond to the simple setting of generalization to states with different support and the last two rows correspond to non-overlapping support. Numbers in the parenthesis represent the standard error over 3 seeds. The best performing method is labeled with blue color. The 2 PPO + 1 Rand column shows the effect of replacing one PPO trained agent with a random agent as a baseline for comparison against the 2 PPO + 1 Diffusion case.
@@ 4 PPO + 1 Rand 5 PPO BC 4 PPO + 1 Diffusion
Table 10: Success rate of each agent in 100 rounds. The first two rows correspond to the simple setting of generalization to states with different support and the last two rows correspond to non-overlapping support. Numbers in the parenthesis represent the standard error over 3 seeds. The best performing method is labeled with blue color. The 4 PPO + 1 Rand column shows the effect of replacing one PPO trained agent with a random agent as a baseline for comparison against the 4 PPO + 1 Diffusion case.
C.8.4 Detailed Results for Ablation
@@
[…]
C.8.5 2v2
The success rates for StarCraft 2v2 are shown below in Table 13. We can see that out-of-combination cases did not cause the performance to drop drastically for MAPPO. This is because the number of combinations in 2v2 is very limited (e.g. $a$,$b$,abaa,$b$,ab), and if one agent dies, MAPPO has encountered scenarios of playing with each unit type individually, therefore falling back to in distribution state again. This scenario also exists for 5v5 and 3v3 but only at the end of each game when only one agent is left.
@@
[…]
C.8.6 More Rendering of States Predicted by the Diffusion model
@@
C.9 Model Runtime and GPU Memory
We include the training time and GPU memory used for the conditioned diffusion model below in Table 14 and 15.
@@ C.9 Model Runtime and GPU Memory
C.9.1 Substituting More MAPPO Agents with Diffusion Agents